OpenStudy (anonymous):

write a polynomial equation in standard form the has the solutions -4 and 5. Help, I do not understand.

OpenStudy (mertsj):

If 5 is an answer then (x-5)=0 If -4 is an answer then (x-(-4))=0 So the equation must be (x-5)(x+4)=0 or x^2-x-20)=0

OpenStudy (anonymous):

thank you. I understand now.

OpenStudy (mertsj):

yw